OpenStudy (anonymous):

What is the fourth term in the expansion of (x + 2)6?

OpenStudy (waleed_imtiaz):

U have to apply binomial theorem to expand this..... do u know binomial theorem ?

OpenStudy (anonymous):

no :/ but my options are: 480x 168x^4 160x^3 120x^3

OpenStudy (amistre64):

options are pointless without knowing how to pick one

OpenStudy (amistre64):

the powers descend from: 6 5 4 3, so 4th term needs to be a power of 3

OpenStudy (anonymous):

oh, so it has to be either C or D?

OpenStudy (amistre64):

that does narrow the options yes

OpenStudy (waleed_imtiaz):

Ya. right @amistre64 if u don't know how to expand , then options are useless....But in my view.. the fourth term will be 40x^3..... can anybody tel me where I made the mistake.... :(

OpenStudy (anonymous):

I'm thinking it's D....?

OpenStudy (amistre64):

i like to set up the general stuff like this; using the power, lets layout the first and last terms 1 2 3 4 5 6 7 term x^6 x^5 x^4 x^3 x^2 x^1 x^0 firsts 2^0 2^1 2^2 2^3 2^4 2^5 2^6 lasts so this thing is at least 2^4 x^3 but there is a leading coeff that would have to be worked out

OpenStudy (amistre64):

6C0 6C1 6C2 6C3 <--- thats the one

OpenStudy (amistre64):

6.5.4 ---- = 20 3.2 20*2^4*x^3

OpenStudy (anonymous):

so it is D? :)

OpenStudy (amistre64):

2^3 .... 8*20 = 160

OpenStudy (amistre64):

waleed; 20*2 was what you did, instead of 20*2^3

OpenStudy (amistre64):

1 6 15 20 15 6 1 coeffs x^6 x^5 x^4 x^3 x^2 x^1 x^0 firsts 2^0 2^1 2^2 2^3 2^4 2^5 2^6 lasts

OpenStudy (anonymous):

ohh, hrm, so it's 160x^3, right

OpenStudy (amistre64):

yes :)

OpenStudy (waleed_imtiaz):

I remember the binomial theorem as this... and it is the transformed form of the theorem without combinations..... it is given below...... (a+b)^n =a^n +a^n-1(b)+n(n-1)/2{a^n-2(b)}^2+n(n-1)(n-2)/6{a^n-3(b)^3}...............+b^n so this is the fourth term..n(n-1)(n-2)/6{a^n-3(b)^3} right put a=x and b=2 and n=6.....

OpenStudy (amistre64):

yes

OpenStudy (waleed_imtiaz):

ya. i got it...... 160x^3 is the answer ..I did wrong calculation last time but the formula was the sam.e

OpenStudy (anonymous):

thank you!

OpenStudy (waleed_imtiaz):

My pleasure......